OpenStudy (anonymous):

Math Analysis: Solve for 0

OpenStudy (anonymous):

treat it as a quadratic equation. let sinx = t so we have t^2 + t -1 =0 t = (-1 + sqrt5)/2 and t= (-1 - sqrt5)/2

OpenStudy (anonymous):

where did you get sqrt5

OpenStudy (anonymous):

b^2 -4ac b=1 , a=1 , c= -1

OpenStudy (anonymous):

i still dont understand. :|

OpenStudy (anonymous):

what's the quadratic equation again?

OpenStudy (anonymous):

could you work out the equation in the quadratic formula that got ou t=-1+sqrt5/2 and t=-1-sqrt5/2

OpenStudy (anonymous):

t^2 + t -1 = 0 t = (-b + sqrt D)/2a and t = ( -b - sqrt D)/2a b = 1 , a = 1 and D = b^2 - 4ac , where c = -1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how did you get the answers

OpenStudy (anonymous):

what do you mean? how did you get the answers? or what did you get as answers?

OpenStudy (anonymous):

i got t=4 for both